OpenStudy (anonymous):

f(2)=6, f'(2)=-4, and f''(2)=2. What is d^2(f^3(x))/dx^2 at x=2? PLEASE EXPLAIN!

OpenStudy (amistre64):

f(2)=6 f^3(2) = 6^3 = 216

OpenStudy (amistre64):

the position at x=2 is 6 the slope at x=2 is -4 the concavity at x=2 is cave up is what i think I can parse

myininaya (myininaya):

I have this one somewhere

OpenStudy (amistre64):

whew!! good lol

OpenStudy (anonymous):

thank you!